OpenStudy (anonymous):

In need of Help !!! Wanted:Someone who is good at using distributive property to rewrite equations And need this person ASAP Reward- Having one extra medal AND fan So please HELP!!! ^_~

OpenStudy (anonymous):

omg i love ur nail pic

OpenStudy (anonymous):

thx my aunt is a nail artist and she lives with me so I get my nails done for free plus she does some really cool things

OpenStudy (anonymous):

So can someone help me ?

OpenStudy (anonymous):

oh cool i love nail art

OpenStudy (anonymous):

ok what is the question

OpenStudy (anonymous):

this picture ^_^

OpenStudy (anonymous):

hold on iam working on it

OpenStudy (anonymous):

okay :)

OpenStudy (anonymous):

Hello?

OpenStudy (anonymous):

@marylou004 can you help or @paki or possibly @Rainbow_dashie

OpenStudy (anonymous):

Hai dashie ^_^

OpenStudy (nirmalnema):

peri =2(l+b)

OpenStudy (nirmalnema):

now by using distri, prop. 2l+2b=peri =2*12.2 +2*9.8=24.4+19.6=44

OpenStudy (nirmalnema):

@!Call~Me~Addi! do u understand

OpenStudy (anonymous):

dashie can you give @ravenwhite799 a medal so I can give @nirmalnema one plz ? and yes @nirmalnema I do understand thanks 4 the help :)

OpenStudy (anonymous):

@Rainbow_dashie do you think you can do that plz :)

OpenStudy (nirmalnema):

ok.. :)

OpenStudy (rainbow_dashie):

imsorry

OpenStudy (anonymous):

ok so i just finished and ur question and i got 44 sorry is took so long i did it the long way

OpenStudy (anonymous):

what ?

OpenStudy (rainbow_dashie):

huh

OpenStudy (nirmalnema):

thanks for the medal...!!

OpenStudy (anonymous):

your welcome @nirmalnema and @ravenwhite799 give my trusty friend @Rainbow_dashie a medal plz :)

OpenStudy (anonymous):

oh... dashie sorry i was talking to @ravenwhite799 because her and @nirmalnema got two VERY diffrent answers

OpenStudy (anonymous):

so i'm wondering which one is right

OpenStudy (nirmalnema):

44 is correct

OpenStudy (anonymous):

for which box there is 2 ?

OpenStudy (nirmalnema):

what ??

OpenStudy (anonymous):

look at the picture there is TWO boxes...

OpenStudy (anonymous):

need the picture again ?

OpenStudy (nirmalnema):

neither box... 1) 12.2 2) 44

OpenStudy (anonymous):

okay that what I needed thank you :)

OpenStudy (anonymous):

@nirmalnema and @ravenwhite799 plz fan me :)

OpenStudy (anonymous):

I ALL READY DID I CAN U PLEASE FAN ME

OpenStudy (anonymous):

i already did ^_^